# D-52056 Aachen, Germany

### v2

@@ 1 Introduction
[…]
@@
On the theory side, NLO QCD corrections for the $pp\to t\bar{t}t\bar{t}$ process with stable top quarks have been calculated for the first time a while ago Bevilacqua:2012em and a few years later recomputed in Refs. Alwall:2014hca ; Maltoni:2015ena . Besides NLO QCD corrections, a further step towards more precise predictions for this process has been achieved by evaluating the complete-NLO corrections Frederix:2017wme that additionally include NLO EW higher-order effects and subleading contributions at perturbative orders from ${\cal O}(\alpha_{s}^{5})$ to ${\cal O}(\alpha^{5})$. However, it has been shown that there are cancellations between different terms in $\alpha_{s}$ and $\alpha$, so in general the higher-order effects for the $pp\to t\bar{t}t\bar{t}$ process are dominated by $\alpha_{s}$ corrections to the leading QCD process at ${\cal O}(\alpha_{s}^{4})$. Finally, very recently the calculation for $pp\to t\bar{t}t\bar{t}$ at the next-to-leading logarithmic accuracy has been carried out vanBeekveld:2022hty including also constant ${\cal O}(\alpha_{s})$ non-logarithmic contributions that do not vanish at the absolute production threshold defined as $M^{2}/\hat{s}\rightarrow 1$, where $M=4t$. In all these studies only the magnitude of the higher-order effects in the total production rate of four top quarks has been investigated, while the reliable description of the fiducial cross sections is still lacking. For realistic studies, the final state with $12$ particles must be considered that consists of four $b$-jets, additional light jets and/or charged leptons together with the (significant) missing transverse momentum. Here, too, reliable theoretical predictions are needed. The first attempt in this direction has already been accomplished in Ref. Jezo:2021smh by matching theoretical predictions for the $pp\to t\bar{t}t\bar{t}$ process with the dipole-style $p_{T}$-ordered parton shower from the Pythia8 Monte Carlo (MC) program Sjostrand:2006za ; Sjostrand:2014zea using the Powheg framework Nason:2004rx ; Alioli:2010xd . In this study, top-quark decays in the $\ell+jets$ channel were modelled at the LO level, preserving spin correlation effects. For the first time, not only the impact of NLO QCD corrections and $t\bar{t}$ spin correlation effects, but also the importance of the subleading EW production modes, that have been included with the LO accuracy only, has been investigated at the differential fiducial cross-section level. In addition, a comparison with the results obtained within the ${}_{-}$aMC@NLO framework Alwall:2014hca , that employs the MC@NLO matching to parton shower programs Frixione:2002ik ; Frixione:2003ei , has been performed. Basically, parton showers represent higher-order corrections to the hard process. However, in this approach an approximation scheme is used in which only colinear parton splitting or soft gluon emission contributions are included. Thus, to cover the entire fiducial phase space it is important to incorporate higher-order effects in top-quark decays already at the matrix element level. This would represent a further step towards the complete description of the jet radiation pattern for this process.
@@ 2 Description of the calculation
Figure 1: Representative Feynman diagrams for the process $pp\to t\bar{t}t\bar{t}\to W^{+}W^{-}W^{+}W^{-}b\bar{b}\,b\bar{b}\to\ell^{+}\nu_{\ell}\,\ell^{-}\bar{\nu}_{\ell}\,\ell^{+}\nu_{\ell}\,\ell^{-}\bar{\nu}_{\ell}\,b\bar{b}\,b\bar{b}$ at the order of $\mathcal{O}(\alpha_{s}^{4}\alpha_{s}^{8})$. For simplicity, the decays of top quarks are not shown in the diagrams. All graphs were generated with FeynGame Harlander:2020cyh .
@@
We have conducted multiple cross-checks to ensure the accuracy of our results. To begin with, we checked the coefficients of the poles in $\epsilon$, i.e. $1/\epsilon$ and $1/\epsilon^{2}$, of the virtual amplitudes for $gg$ and $q\bar{q}$ subprocesses with ${}_{-}$aMC@NLO Alwall:2014hca for a few random phase-space points for stable $t\bar{t}t\bar{t}$ production and for the $t\to W^{+}b\to\ell^{+}\nu_{\ell}\,b$ decay. In the next step, we repeated this procedure using the Recola MC program Actis:2016mpe . This time, however, also the finite parts along with the coefficients of the poles in $\epsilon$ were cross-checked without separating the production stage from the top-quark decays. Because Recola is able to provide matrix elements in the so-called double-pole approximation Denner:2000bj ; Accomando:2004de ; Denner:2016jyo , it was rather straightforward to prepare this program for the use in the NWA by interfacing it to the Helac-NLO system. This approach has already been tested before in the case of NLO QCD calculations for $pp\to t\bar{t}jj$ Bevilacqua:2022ozv and $pp\to t\bar{t}\gamma\gamma$ Stremmer:2023kcd . For the real corrections, we have checked for several phase-space points that the poles in $\epsilon$ that appear in the ${\cal I}$-operator and those appearing in the virtual part of our calculations cancel. We have confirmed that the real emission part of the calculation is independent of any phase-space restriction imposed by the unphysical $\alpha_{\rm max}$ parameter. This check has been applied to both the Catani-Seymour and Nagy-Soper subtraction schemes. Finally, having two independent subtraction schemes allowed us to cross-check the correctness of the real emission part of the calculation in an even more robust way.
@@ 3 Computational setup
at NLO in QCD, where $\ell^{\pm}$ stands for ±$\ell^{\pm}=e^{\pm},\mu^{\pm}$. To perform this calculation we use the NWA. For simplicity, we refer to this process as $pp\to t\bar{t}t\bar{t}$ in the $4\ell$ top-quark decay channel. We do not consider $\tau$ leptons as they are often analysed separately at the LHC due to their rather rich and complex decay pattern. Our results are obtained for LHC Run III center-of-mass energy of $\sqrt{s}=13.6$ TeV. The numerical values we use for the SM input parameters are given below:
$\begin{array}[]{lll}G_{\mu}=1.1663787\cdot 10^{-5}~{}{\rm GeV}^{-2}\,,&&m_{t}=172.5~{}{\rm GeV}\,,\\ m_{W}=80.379~{}{\rm GeV}\,,&&\Gamma_{W}^{\rm NLO}=2.0972~{}{\rm GeV}\,,\\ m_{Z}=91.1876~{}{\rm GeV}\,,&&m_{b}=0~{}{\rm GeV}\,.\end{array}$ (3)
@@ 5 Integrated fiducial cross sections
Table 1: Integrated fiducial cross sections at LO and NLO in QCD for the $pp\to t\bar{t}t\bar{t}$ process in the $4\ell$ channel at the LHC with $\sqrt{s}=$ 13.6 TeV. Results are presented for MSHT20, NNPDF3.1 and CT18 PDF sets. They are evaluated using $\mu_{0}=2m_{t}$ and $\mu_{0}=E_{T}/4$. Also given are theoretical uncertainties coming from the scale variation and from PDFs. In the last column the ${\cal K}$-factor is shown.
@@
[…]
Within a given PDF set we can not assess the differences that enter into the parameterisation of the various PDF sets. To investigate this issue, following the recommendation of the PDF4LHC group for the usage of PDFs suitable for applications at the LHC that are connected to SM processes, we additionally present our results for NNPDF3.1 and CT18 PDF sets. The LO and NLO integrated fiducial cross sections for the three PDF sets are shown in Table 1 for two different scale choices: $\mu_{0}=2m_{t}$ and $\mu_{0}=E_{T}/4$. The $\mathcal{K}$ factor, defined as ${\cal K}=\sigma^{\rm NLO}/\sigma^{\rm LO}$, is also provided in the last column. The $68\%$ C.L. PDF uncertainties for NNPDF3.1 are up to $2\%$ only, whereas for the CT18 PDF set they are of the order of $6\%$. As expected, we observe significant variations in the values of the LO cross section for the three LO PDF sets that are provided for $\alpha_{s}(m_{Z})=0.130$. These differences result in a variety of the $\mathcal{K}$ factor: ${\cal K}=(1.25-1.30)$ for NNPDF3.1 and ${\cal K}=(1.03-1.04)$ for CT18. Using NLO PDF sets instead for the LO calculations would correspond to lower values of $\sigma^{\rm LO}$ and, consequently, higher ${\cal K}$ factors. On the other hand, at the NLO QCD level and independent of the scale setting, the differences among $\sigma^{\rm NLO}$ for the three PDF sets are of the order of $1\%$ only.
@@
Table 2: Integrated fiducial cross sections at NLO in QCD for the $pp\to t\bar{t}t\bar{t}$ process in the $4\ell$ channel at the LHC with $\sqrt{s}=$ 13.6 TeV for three different scenarios: $\sigma^{\rm NLO}_{\rm full}$, $\sigma^{\rm NLO}_{\rm LO_{dec}}$ and $\sigma^{\rm NLO}_{\rm exp}$. All results are given for our default setup with the MSHT20 PDF set and for two different scale choices: $\mu_{0}=2m_{t}$ and $\mu_{0}=E_{T}/4$. Also given are theoretical uncertainties coming from the scale variation. In the last column the ratio to the $\sigma^{\rm NLO}_{\rm exp}$ result is provided.
@@ 6 Differential fiducial cross-section distributions
While interesting, the NLO QCD corrections to the integrated fiducial cross section do not give us a complete picture of the impact of the higher-order effects on the $pp\to t\bar{t}t\bar{t}$ process in the $4\ell$ channel. To understand how NLO QCD corrections affect specific observables and/or phase-space regions, it is necessary to analyse the results at the differential fiducial cross-section level. To this end, we start with the comparison between LO and NLO QCD computations for several observables. Additionally, we can inspect the size of theoretical uncertainties in different parts of the phase space. We present our findings only for the dynamical scale choice $\mu_{R}=\mu_{F}=\mu_{0}=E_{T}/4$ and for the default (N)LO MSHT20 PDF set. It is a well known fact that the fixed scale choice is not adequate at the differential cross-section level. For example, for some values of $\mu_{R}=\mu_{F}=\xi\mu_{0}$, where $\mu_{0}=2m_{t}$, the NLO results in the high-energy tails of dimensionful distributions become negative. Furthermore, in similar phase-space regions the LO and NLO scale variation bands no longer overlap. Finally, it can even happen that the size of the scale variation at the NLO level actually exceeds the one of the LO prediction. All these effects have already been observed and discussed in detail in our previous studies for the $pp\to t\bar{t}+X$ process, where $X=j,\gamma,Z,W^{\pm},H,jj,\gamma\gamma,b\bar{b}$, see e.g. Refs. Bevilacqua:2015qha ; Bevilacqua:2016jfk ; Bevilacqua:2018woc ; Bevilacqua:2019cvp ; Bevilacqua:2020pzy ; Bevilacqua:2021cit ; Stremmer:2021bnk ; Bevilacqua:2022ozv ; Bevilacqua:2022nrm ; Stremmer:2023kcd . Similar effects with a fixed scale setting have been examined for the $pp\to t\bar{t}$ process at NLO in QCD in e.g. Ref. Denner:2012yc .

### v3

@@ 1 Introduction
With its short lifetime, however, the top quark decays before top-flavoured hadrons or $t\bar{t}$-bound states can be formed. We therefore only have access to the properties of the top quark through its decay products. Given that the top quark decays almost $100\%$ of the time into $t\to Wb$, its final states are classified via the $W$ gauge boson decays into $W\to\ell\nu_{\ell}$ and $W\to q\bar{q}^{\,\prime}$. Consequently, the SM $pp\to t\bar{t}t\bar{t}\to W^{+}W^{-}W^{+}W^{-}b\bar{b}b\bar{b}$ process produces a plethora of final states, all of which yield interesting and distinctive signatures at the LHC. The dominant $4t$ decay mode is the mono-leptonic channel with a branching ratio $({\cal BR})$ of $40\%$. It is followed by the fully hadronic and opposite-sign dilepton modes with ${\cal BR}\approx 20\%$ each, the same-sign dilepton and the trilepton modes with ${\cal BR}\approx 10\%$ each and finally by the fully leptonic mode with the smallest contribution of the order of $1.2\%$ Alvarez:2016nrz . The last branching ratio is further reduced to $0.2\%$ when ±$\ell^{\pm}=e^{\pm},\mu^{\pm}$ are only considered and the $\tau$ lepton is not taken into account. Both ATLAS and CMS have recently observed the production of four top-quarks with the LHC Run II energy of $13$ TeV using the integrated luminosity of about $140$ fb-1 ATLAS:2023ajo ; CMS:2023ftu . Events with two same-sign, three, or four charged leptons (electrons and muons) and additional light- and/or $b$-jets have been analysed to achieve the required significance of more than $5\sigma$ that is needed to claim observation. Independent of the experiment and taking into account current uncertainties of the measurements, the obtained $t\bar{t}t\bar{t}$ cross sections are in agreement with the SM prediction. At present, however, both ATLAS and CMS measurements are limited by statistical uncertainties. The largest systematic uncertainties, on the other hand, come from the modelling of the additional jet activity in $t\bar{t}W^{\pm}$ production, the normalisation of the $pp\to t\bar{t}Z$ background process and the modelling of the $pp\to t\bar{t}t\bar{t}$ signal process. Therefore, an accurate theoretical description of $t\bar{t}t\bar{t}$ production at the LHC comprising all the features of the process is much needed and more relevant than ever.
@@
On the theory side, NLO QCD corrections for the $pp\to t\bar{t}t\bar{t}$ process with stable top quarks have been calculated for the first time a while ago Bevilacqua:2012em and a few years later recomputed in Refs. Alwall:2014hca ; Maltoni:2015ena . Besides NLO QCD corrections, a further step towards more precise predictions for this process has been achieved by evaluating the complete-NLO corrections Frederix:2017wme that additionally include NLO EW higher-order effects and subleading contributions at perturbative orders from ${\cal O}(\alpha_{s}^{5})$ to ${\cal O}(\alpha^{5})$. However, it has been shown that there are cancellations between different terms in $\alpha_{s}$ and $\alpha$, so in general the higher-order effects for the $pp\to t\bar{t}t\bar{t}$ process are dominated by $\alpha_{s}$ corrections to the leading QCD process at ${\cal O}(\alpha_{s}^{4})$. Finally, very recently the calculation for $pp\to t\bar{t}t\bar{t}$ at the next-to-leading logarithmic accuracy has been carried out vanBeekveld:2022hty including also constant ${\cal O}(\alpha_{s})$ non-logarithmic contributions that do not vanish at the absolute production threshold defined as $M^{2}/\hat{s}\rightarrow 1$, where $M=4t$. In all these studies only the magnitude of the higher-order effects in the total production rate of four top quarks has been investigated, while the reliable description of the fiducial cross sections is still lacking. For realistic studies, the final state with $12$ particles must be considered that consists of four $b$-jets, additional light jets and/or charged leptons together with the (significant) missing transverse momentum. Here, too, reliable theoretical predictions are needed. The first attempt in this direction has already been accomplished in Ref. Jezo:2021smh by matching theoretical predictions for the $pp\to t\bar{t}t\bar{t}$ process with the dipole-style $p_{T}$-ordered parton shower from the Pythia8 Monte Carlo (MC) program Sjostrand:2006za ; Sjostrand:2014zea using the Powheg framework Nason:2004rx ; Alioli:2010xd . In this study, top-quark decays in the $\ell+jets$ channel have been modelled at LO. Consequently, spin correlations are also described at the LO level in the perturbative expansion. For the first time, not only the impact of NLO QCD corrections and $t\bar{t}$ spin correlation effects, but also the importance of the subleading EW production modes, that have been included with the LO accuracy only, has been investigated at the differential fiducial cross-section level. In addition, a comparison with the results obtained within the MG5-aMC@NLO framework Alwall:2014hca , that employs the MC@NLO matching to parton shower programs Frixione:2002ik ; Frixione:2003ei , has been performed. Basically, parton showers represent higher-order corrections to the hard process. However, in this approach an approximation scheme is used in which only collinear parton splitting or soft gluon emission contributions are included. Thus, to cover the entire fiducial phase space it is important to incorporate higher-order effects in top-quark decays already at the matrix element level. This would represent a further step towards the complete description of the jet radiation pattern for this process.
@@ 2 Description of the calculation
Figure 1: Representative Feynman diagrams for the process $pp\to t\bar{t}t\bar{t}\to W^{+}W^{-}W^{+}W^{-}b\bar{b}\,b\bar{b}\to\ell^{+}\nu_{\ell}\,\ell^{-}\bar{\nu}_{\ell}\,\ell^{+}\nu_{\ell}\,\ell^{-}\bar{\nu}_{\ell}\,b\bar{b}\,b\bar{b}$ at the order of $\mathcal{O}(\alpha_{s}^{4}\alpha^{8})$. For simplicity, the decays of top quarks are not shown in the diagrams. All graphs were generated with FeynGame Harlander:2020cyh .
@@
We have conducted multiple cross-checks to ensure the accuracy of our results. To begin with, we checked the coefficients of the poles in $\epsilon$, i.e. $1/\epsilon$ and $1/\epsilon^{2}$, of the virtual amplitudes for $gg$ and $q\bar{q}$ subprocesses with MadGraph-aMC@NLO Alwall:2014hca for a few random phase-space points for stable $t\bar{t}t\bar{t}$ production and for the $t\to W^{+}b\to\ell^{+}\nu_{\ell}\,b$ decay. In the next step, we repeated this procedure using the Recola MC program Actis:2016mpe . This time, however, also the finite parts along with the coefficients of the poles in $\epsilon$ were cross-checked without separating the production stage from the top-quark decays. Because Recola is able to provide matrix elements in the so-called double-pole approximation Denner:2000bj ; Accomando:2004de ; Denner:2016jyo , it was rather straightforward to prepare this program for the use in the NWA by interfacing it to the Helac-NLO system. This approach has already been tested before in the case of NLO QCD calculations for $pp\to t\bar{t}jj$ Bevilacqua:2022ozv and $pp\to t\bar{t}\gamma\gamma$ Stremmer:2023kcd . For the real corrections, we have checked for several phase-space points that the poles in $\epsilon$ that appear in the ${\cal I}$-operator and those appearing in the virtual part of our calculations cancel. We have confirmed that the real emission part of the calculation is independent of any phase-space restriction imposed by the unphysical $\alpha_{\rm max}$ parameter. This check has been applied to both the Catani-Seymour and Nagy-Soper subtraction schemes. Finally, having two independent subtraction schemes allowed us to cross-check the correctness of the real emission part of the calculation in an even more robust way.
@@ 3 Computational setup
at NLO in QCD, where $\ell^{\pm}$ stands for ±$\ell^{\pm}=e^{\pm},\mu^{\pm}$. If instead $\ell=e^{\pm},\mu^{\pm},\tau^{\pm}$, then all the results presented in this paper should be multiplied by the following factor $81/16\approx 5.0625$ to obtain the correct normalisation. To perform this calculation we use the NWA. For simplicity, we refer to this process as $pp\to t\bar{t}t\bar{t}$ in the $4\ell$ top-quark decay channel. We do not consider $\tau$ leptons as they are often analysed separately at the LHC due to their rather rich and complex decay pattern. Our results are obtained for LHC Run III center-of-mass energy of $\sqrt{s}=13.6$ TeV. The numerical values we use for the SM input parameters are given below:
$\begin{array}[]{lll}G_{\mu}=1.1663787\cdot 10^{-5}~{}{\rm GeV}^{-2}\,,&&m_{t}=172.5~{}{\rm GeV}\,,\vspace{0.2cm}\\ m_{W}=80.379~{}{\rm GeV}\,,&&\Gamma_{W}^{\rm NLO}=2.0972~{}{\rm GeV}\,,\vspace{0.2cm}\\ m_{Z}=91.1876~{}{\rm GeV}\,,&&m_{b}=0~{}{\rm GeV}\,.\end{array}$ (3)
@@ 5 Integrated fiducial cross sections
Table 1: Integrated fiducial cross sections at LO and NLO in QCD for the $pp\to t\bar{t}t\bar{t}$ process in the $4\ell$ channel at the LHC with $\sqrt{s}=$ 13.6 TeV. Results are presented for MSHT20, NNPDF3.1 and CT18 PDF sets. They are evaluated using $\mu_{0}=2m_{t}$ and $\mu_{0}=E_{T}/4$. Also given are theoretical uncertainties coming from the scale variation and from PDFs. MC integration errors are displayed in parentheses. In the last column the ${\cal K}$-factor is shown.
Scale Order PDF $\sigma$ [ab] $\mathcal{K}=\sigma^{\textrm{NLO}}/\sigma^{\textrm{LO}}$
$\mu_{0}=2m_{t}$ LO NNPDF3.1_lo_as_0130 $3.7389(2)^{+72\%}_{-39\%}$ 1.30
LO NNPDF3.1_lo_as_0118 2.8835($2.8835(2)^{+65\%}_{-37\%}$ 1.68
LO NNPDF3.1_nlo_as_0118 $3.1068(2)^{+68\%}_{-38\%}$ 1.56
NLO NNPDF3.1_nlo_as_0118 $4.846(2)^{+13\%}_{-21\%}$
$\mu_{0}=E_{T}/4$ LO NNPDF3.1_lo_as_0130 $4.0930(3)^{+73\%}_{-39\%}$ 1.25
LO NNPDF3.1_lo_as_0118 3.1356($3.1356(3)^{+66\%}_{-37\%}$ 1.64
LO NNPDF3.1_nlo_as_0118 $3.3633(2)^{+69\%}_{-38\%}$ 1.52
NLO NNPDF3.1_nlo_as_0118 $5.126(3)^{+12\%}_{-21\%}$
Table 2: Integrated fiducial cross sections at LO and NLO in QCD for the $pp\to t\bar{t}t\bar{t}$ process in the $4\ell$ channel at the LHC with $\sqrt{s}=$ 13.6 TeV. The LO and NLO NNPDF3.1 PDF sets are employed for LO predictions with different values of $\alpha_{s}(m_{Z})$ using $\mu_{R}=\mu_{F}=\mu_{0}$ where $\mu_{0}=2m_{t}$ and $\mu_{0}=E_{T}/4$. Also given are NLO predictions and theoretical uncertainties coming from the scale variation. MC integration errors are displayed in parentheses. In the last column the ${\cal K}$-factor is also shown.
@@
Figure 4: Contour plots for the scale dependence in the ($\mu_{R},\mu_{F}$) plane at LO (upper plots) and NLO (lower plots) for our default PDF set MSHT20 and two scale choices $\mu_{0}=2m_{t}\;\textrm{and}\;\mu_{0}=E_{T}/4$. Results are presented for the $pp\to t\bar{t}t\bar{t}$ process in the $4\ell$ channel at the LHC with $\sqrt{s}=$ 13.6 TeV.
Within a given PDF set we can not assess the differences that enter into the parameterisation of the various PDF sets. To investigate this issue, following the recommendation of the PDF4LHC group for the usage of PDFs suitable for applications at the LHC that are connected to SM processes, we additionally present our results for NNPDF3.1 and CT18 PDF sets. The LO and NLO integrated fiducial cross sections for the three PDF sets are shown in Table 1 for two different scale choices: $\mu_{0}=2m_{t}$ and $\mu_{0}=E_{T}/4$. The $\mathcal{K}$ factor, defined as ${\cal K}=\sigma^{\rm NLO}/\sigma^{\rm LO}$, is also provided in the last column. The $68\%$ C.L. PDF uncertainties for NNPDF3.1 are up to $2\%$ only, whereas for the CT18 PDF set they are of the order of $6\%$. We observe substantial variations in the values of the LO cross section for the three LO PDF sets that are provided for $\alpha_{s}(m_{Z})=0.130$. These differences, ranging from $5\%$ to $15\%$, result in a variety of the $\mathcal{K}$ factor: ${\cal K}=(1.25-1.30)$ for NNPDF3.1 and ${\cal K}=(1.03-1.04)$ for CT18. As expected, using NLO PDF sets instead for the LO calculations would correspond to lower values of $\sigma^{\rm LO}$ and, consequently, higher ${\cal K}$ factors. Similar effects would be observed when using LO PDF sets with $\alpha_{s}(m_{Z})=0.118$. In order to substantiate these statements in Table 2 we present LO integrated fiducial cross sections as obtained with the following PDF sets provided by the NNPDF collaboration: the LO NNPDF3.1 PDF set with $\alpha_{s}(m_{Z})=0.130$ and $\alpha_{s}(m_{Z})=0.118$ as well as the NLO NNPDF3.1 PDF set. For consistency and readability, the NLO cross section is also displayed. Again, the results are given for two scale settings $\mu_{0}=2m_{t}$ and $\mu_{0}=E_{T}/4$. We can observe that the LO results are simply not adequate to describe the $pp\to t\bar{t}t\bar{t}$ process in the $4\ell$ channel. They are sensitive not only to scale variation, but also to the choice of PDF set and the value of $\alpha_{s}(m_{Z})$ used. This large dependence is reflected in the obtained ${\cal K}$-factor, which for the process at hand ranges from ${\cal K}=1.25$ to ${\cal K}=1.68$ within a given PDF set and from ${\cal K}=1.03$ to ${\cal K}=1.68$ once the three PDF sets, which are recommended by the PDF4LHC working group, are employed. On the other hand, at the NLO QCD level and independent of the scale setting, the differences among $\sigma^{\rm NLO}$ for the three PDF sets (MSHT20, NNPDF3.1 and CT18) are of the order of $1\%$ only. Generally, as explained by the various PDF collaborations there are large differences between the LO and NLO PDF sets, both for central values and uncertainties. In particular, the LO uncertainties are significant due to the poor quality of the fit. This causes the shift in quark PDFs that can be substantial, while the gluon PDF at small $x$ is completely different between LO and NLO, see e.g. Ref. NNPDF:2017mvq . This is due to the fact that the singular small-$x$ behaviour of the quark to gluon splittings only starts at NLO. In addition, the PDF global analyses involve data from deep-inelastic scattering (DIS) experiments, fixed-target Drell-Yan (DY) data, and collider measurements from the Tevatron and LHC. However, gluon initiated DIS and DY processes, e.g. electroweak boson production processes from ATLAS, CMS and LHCb that are very important for the determination of the gluon PDF, vanish at LO.
@@
Table 3: Integrated fiducial cross sections at NLO in QCD for the $pp\to t\bar{t}t\bar{t}$ process in the $4\ell$ channel at the LHC with $\sqrt{s}=$ 13.6 TeV for three different scenarios: $\sigma^{\rm NLO}_{\rm full}$, $\sigma^{\rm NLO}_{\rm LO_{dec}}$ and $\sigma^{\rm NLO}_{\rm exp}$. All results are given for our default setup with the MSHT20 PDF set and for two different scale choices: $\mu_{0}=2m_{t}$ and $\mu_{0}=E_{T}/4$. Also given are theoretical uncertainties coming from the scale variation. MC integration errors are displayed in parentheses. In the last column the ratio to the $\sigma^{\rm NLO}_{\rm exp}$ result is provided.
@@ 6 Differential fiducial cross-section distributions
While interesting, the NLO QCD corrections to the integrated fiducial cross section do not give us a complete picture of the impact of the higher-order effects on the $pp\to t\bar{t}t\bar{t}$ process in the $4\ell$ channel. To understand how NLO QCD corrections affect specific observables and/or phase-space regions, it is necessary to analyse the results at the differential fiducial cross-section level. To this end, we start with the comparison between LO and NLO QCD computations for several observables. Additionally, we can inspect the size of theoretical uncertainties in different parts of the phase space.
It should be mentioned here that measuring the differential cross-section distributions shown below will be rather difficult even at the High-Luminosity phase of the LHC. The reason behind this is the expected number of events for this process. Indeed, considering only ±$\ell^{\pm}=e^{\pm},\mu^{\pm}$, the expected number of events is $N=(15-20)$ per experiment, assuming an integrated luminosity of $(3-4)~{}{\rm ab}^{-1}$. On the other hand, when ±$\ell^{\pm}=e^{\pm},\mu^{\pm},\tau^{\pm}$ we can estimate $N=(157-209)$ events for both ATLAS and CMS. We note that the numbers given here correspond to our definition of the fiducial phase-space region. These numbers will change when more inclusive event selection cuts are applied. Nevertheless, proper modelling of differential cross-section distributions is important for example for the many preparatory studies and various analyses that are currently being conducted for this process. In addition, the analysis at the differential cross-section level for the SM $pp\to t\bar{t}t\bar{t}t+X$ process is a necessary step towards a correct interpretation of possible signals of new physics that may arise in the $4\ell$ channel.
@@
We present our findings only for the dynamical scale choice $\mu_{R}=\mu_{F}=\mu_{0}=E_{T}/4$ and for the default (N)LO MSHT20 PDF set. It is a well known fact that the fixed scale choice is not adequate at the differential cross-section level. For example, for some values of $\mu_{R}=\mu_{F}=\xi\mu_{0}$, where $\mu_{0}=2m_{t}$, the NLO results in the high-energy tails of dimensionful distributions become negative. Furthermore, in similar phase-space regions the LO and NLO scale variation bands no longer overlap. Finally, it can even happen that the size of the scale variation at the NLO level actually exceeds the one of the LO prediction. All these effects have already been observed and discussed in detail in our previous studies for the $pp\to t\bar{t}+X$ process, where $X=j,\gamma,Z,W^{\pm},H,jj,\gamma\gamma,b\bar{b}$, see e.g. Refs. Bevilacqua:2015qha ; Bevilacqua:2016jfk ; Bevilacqua:2018woc ; Bevilacqua:2019cvp ; Bevilacqua:2020pzy ; Bevilacqua:2021cit ; Stremmer:2021bnk ; Bevilacqua:2022ozv ; Bevilacqua:2022nrm ; Stremmer:2023kcd . Similar effects with a fixed scale setting have been examined for the $pp\to t\bar{t}$ process at NLO in QCD in e.g. Ref. Denner:2012yc .